Find the power function that the graph of f resembles for large values of |x|.
f(x) = -x^2(x + 4)^3(x^2 - 1)
a. y = x^7
b. y = -x^7
c. y = x^3
d. y = x^2

Answer by jim_thompson5910(35256)

-x^2 is the leading term for the first factor

For the second factor, (x+4)^3 has the leading term x^3

The third factor has the leading term x^2

Multiply the leading terms:
-x^2*x^3*x^2 = -x^(2+3+2) = -x^7

----------------------------------------------------------------

The leading term of f(x) = -x^2(x + 4)^3(x^2 - 1) is -x^7

So the answer is choice B) -x^7